Post Card Depot, an large retailer of post cards, orders 9,123,934 post cards per year from its manufacturer. Post Card Depot plans on ordering post card 8 times over the next year. Post Card Depot receives the same number of post cards each time it orders. The carrying cost is $0.22 per post card per year. The ordering cost is $285 per order.

What is the annual total inventory management costs of post card inventory?
Answer #1

Order size = 9123934/8 = 1,140,491.75

Annual ordering cost = 285*8 = $2,280

Annual carrying cost = Average inventory*Carrying cost per unit per year

= (1,140,491.75+0)/2*0.22

= $125,454.0925

Total annual inventory management cost = Ordering cost + Carrying cost

= $127,734.0925